Preparation of the correct buffer is key to any good biological experiment and it is important that you understand how to calculate the mass of each chemical required to make that buffer and what the resulting concentration of those constituents will be in moles per litre.

Your text book explains that moles are just a way to express the amount of a substance, such that one mole is equal to 6.02 x 1023 particles of that substance. These particles can be can be atoms, molecules, ions etc, so 1 mole of water is equal to 6.02 x 1023 water molecules, or 1 mole of Na+ is equal to 6.02 x 1023  Na+ ions. Since different chemicals have different molecular weights (based on the number of protons and neutrons each atom contains) 1 mole or 6.02 x 1023 atoms of oxygen (O) will have a mass of 16g whereas 1 mole or 6.02 x 1023 atoms of sodium (Na) will have a mass of 23g

If you need more information on moles, please read Encyclopedia Britannica's Moles website. 

Although you may sometimes see it written as g/litre, the concentration of solutions is more often described in term of molarity since it better defines the chemical properties of a solution because it is proportional to the number of molecules or ions in solution, irrespective of molecular mass of its constituents. However, it is not possible to measure moles on a laboratory balance, so in the first instance chemicals are measured by mass (milligrams, grams, kilograms etc) and the number of moles is calculated using the known molecular mass (often called molecular weight and abbreviated to M.W.) of the chemical. As indicated earlier, the molecular mass of a chemical is based on the number of protons and neutrons that is contained in each atom (eg NaCl is made up of one molecule of Na, M.W. = 22.99g and one molecule of Cl, M.W. = 35.45g, so the M.W. of NaCl is 58.44g). These values can be found in the periodic table however the molecular mass of chemicals is generally provided by any vendors of the products and so can also be found on various suppliers’ websites.

 

When the concentrations of solutions are as described as ‘molar’, this refers to number of moles per litre eg a 3-molar solution of NaCl will contain 3 moles of NaCl in 1 litre of water. As indicated above, the M.W. of NaCl is 58.44g, so in 58.44g there are 6.02 x 1023 NaCl molecules ie 1 mole. So, for 3 moles of NaCl you would need to dissolve 175.32g in 1 litre of water (175.32/58.44 =3) whereas If you only dissolved 29.22g of NaCl in 1 litre of water this would result in a 0.5 molar solution (29.22/58.44= 0.5)

 

1.     As directed you need to check the periodic table and pick up the atomic masses for each of the component atoms in the compounds. For example, for NaCl you need to pick the atomic weight of both sodium and chlorine and then add them to two decimal places to obtain the molecular mass of NaCl. Be sure to multiply the atomic masses by the number of individual atoms of the same element present in each compound before finally adding to the masses of other component atoms of other elements to make up the total molecular masses. 

 

2.     From there you can calculate the number of 'moles' of each compound by multiplying the provided weight of compound used in the PBS solution by their respective molar mass conversion factors (i.e. 1L divided by the molecular mass you have calculated in the first step) 

 

3.     Now, the molarity in Mol per Litre (mol/l) is given by the 'number of moles' of each compound (calculated in step 2 above) divided by the given volume of the solution.

For more information on how to calculate morality, refer to wikiHow's 4 Ways to Calculate Molarity.

Using periodic table found in your textbook, calculate (to 2 decimal places) the molecular mass for each of the compounds used to make PBS.

Create the following table and fill it in with the mass of each component required to make 1 litre of 10 x PBS (the recipe for 10x PBS is below question 2) and their final molar concentration in the buffer calculated as described above.

 

Compound formula

 

Molecular mass (in g/mol)

Mass of compound per litre of 10x PBS (in g)

Molar concentration (in mol/l)

NaCl

 

 

KCl

 

 

Na2HPO4

 

 

KH2PO4

 

 


1
Expert's answer
2021-02-10T05:08:04-0500

Create the following table and fill it in with the mass of each component required to make 1 litre of 10Х PBS (the recipe for 10X PBS is below question 2) and their final molar concentration in the buffer calculated as described above.


Solution:

Since there is no recipe of 1 litre of 10X FBS in the task, let`s use the recipe* from company that sells materials, reagents and equipment for the life sciences research: 17.8 g of Na2HPO4 • 2H2O, 2.4 g of KH2PO4, 80 g of NaCl, 2 g of KCl.

* - (n.d.) 10X Phosphate-Buffered Saline (PBS) Recipe Calculator

Retrieved February, 9, 2021, from https://www.sigmaaldrich.com/technical- documents/protocols/biology/western-blotting/buffers-recipes/10x-phosphate-buffered-saline.html


The number of moles of Na2HPO4 • 2H2O is equal to the number of moles of Na2HPO4.

M.W. of Na2HPO4 • 2H2O = 2 • 22.99 + 1 + 30.97 + 4 • 15.99 + 2 • (2 • 1 + 15.99) = 177.99

M.W. of KH2PO4 = 39.098 + 2 • 1 + 30.97 + 4 • 15.99 = 136.09

M.W. of NaCl = 22.99 + 35.45 = 58.44

M.W. of KCl = 39.098 + 35.45 = 74.55

M.W. of Na2HPO4 = 2 • 22.99 + 1 + 30.97 + 4 • 15.99 = 142.01

M.W. 2H2O = 2 • (2 • 1 + 15.99) = 35.98


Number of moles of Na2HPO4 • 2H2O = 17.8/177.99 = 100 • 10-3

N.M. of KH2PO4 = 2.4/136.09 = 18 • 10-3

N.M. of NaCl = 80/58.44 = 1.37

N.M. of KCl = 2/74.55 = 27 • 10-3


Mass of Na2HPO4 = 17.8 g - 100 • 10-3 • 35,98 = 14.2 g


Molar concentration is equal to mole number because volume of solution is 1 litre.


Answer:

NaCl: 58.44 g/mole, 80 g, 1.37 M

KCl: 74.55 g/mole, 2 g, 27 mM

Na2HPO4: 142.01 g/mole, 14.2 g, 100 mM

KH2PO4: 136.09 g/mole, 2.4 g, 18 mM